# CARMA: Enhanced Compositionality in LLMs via Advanced Regularisation and Mutual Information Alignment

### v1

@@ Abstract
Nura Aljaafari1,4†,  Danilo S. Carvalho1,3,  André Freitas1,2,3 1 Department of Computer Science, University of Manchester, United Kingdom 2 Idiap Research Institute, Switzerland 3 National Biomarker Centre, CRUK-MI, Univ. of Manchester, United Kingdom 4 King Faisal University, Al Hofuf, KSA {firstname.lastname}@[postgrad.]manchester.ac.uk
@@ 1 Introduction
Despite their strong performance across various natural language processing tasks, large language models (LLMs) exhibit persistent weaknesses in compositional generalisation Hupkes et al. (2020); Kim and Linzen (2020a); Aljaafari et al. (2024). These limitations stem from multiple factors, including training objectives and model architectures. Standard autoregressive training methods, such as next-token prediction, prioritise statistical correlations in token sequences over structured semantic understanding Yin et al. (2023a); Dziri et al. (2024). As a result, token representations often lack structured compositionality, leading to fragmented information processing within layers (horizontal misalignment) and across layers (vertical inconsistency).
@@ 2 Compositionality in LLMs
Compositional generalisation (CG) in linguistics encompasses five key principles: systematicity, productivity, substitutivity, localism, and over-generalisation Dankers et al. (2022a). These principles have been explored in LLMs for various applications, including compositional instruction Yang et al. (2024b), semantic parsing Li et al. (2023), machine translation Li et al. (2021), and multi-step inference Zhang et al. (2024). Empirical studies reveal that standard Transformer-based LLMs exhibit limited CG, even for relatively simple compositional tasks. For instance, models frequently struggle to assemble tokens into words or construct morphemes into coherent structures Aljaafari et al. (2024); Ismayilzada et al. (2024). These limitations are linked to architectural constraints, training objectives, and tokenisation practices, which fragment information and increase sensitivity to input order and contextual noise Murty et al. (2023).
@@
Architectural Mechanisms and Compositional Consistency. Beyond training objectives, architectural mechanisms such as dropout and self-attention contribute to the dispersion of information across the model. This fragmentation increases sensitivity to input order and context, often resulting in errors that undermine compositional consistency Sajjadi et al. (2016); Cai et al. (2021)—the model’s ability to maintain produce consistent outputs when processing variations of semantically equivalent inputs through transformations like word substitution or paraphrasing.
These challenges impact both high-complexity reasoning tasks and simpler operations that demand consistent morphological and syntactic processing Ismayilzada et al. (2024).
Existing Approaches to Enhance CG in LLMs. To address CG limitations, research has explored architectural adjustments, regularisation techniques, and task-specific strategies. For instance, Ontanon et al. (2022) demonstrated that combining relative positional encoding with embeddings enhances CG, particularly in algorithmic tasks. Their findings suggest that weight sharing and copy decoders help retain input structures, thus improving CG accuracy. Other architectural modifications, such as Pushdown Layers Murty et al. (2023) and GroCoT Sikarwar et al. (2022), incorporate mechanisms for tracking syntactic depth and spatial relations, which enable recursive processing of compositional structures.
Models like RegularGPT Chi et al. (2023) introduce adaptive depth and memory mechanisms to facilitate CG by constructing complex structures from simpler components. Studies by Csordás et al. (2021) and Petty et al. (2024) evaluate model depth, parameter configurations, and encoding methods, revealing that architectural choices and training setups—such as avoiding early stopping and prioritising accuracy over loss minimisation—are critical to enhancing CG. In neural machine translation (NMT), Dankers et al. (2022b) reformulated CG evaluations, finding a positive correlation between data size and compositional performance, underscoring the importance of extensive, real-world benchmarks for capturing the complexities of linguistic compositionality.
Frameworks like CompMCTG and Meta-MCTG Zhong et al. (2024) offer benchmarks for evaluating CG in multi-aspect text generation, suggesting that joint training and meta-learning approaches can improve fluency. However, significant performance drops persist in out-of-distribution tasks. Additionally, synthetic tasks reveal that recursive, step-by-step prompt formats support combinatorial generalisation, although training biases and sequence order constraints remain limiting factors Ramesh et al. (2024).
@@ 3 Enhanced Compositionality via Advanced Regularisation and Mutual Information Alignment (CARMA)
This section formalises compositionality, introduces the core principles of CARMA, and details its components. Figure 1 illustrates the CARMA method, highlighting its optimisation process and key components.
@@ 3.2 CARMA Formalisation
where $f^{(k)}(X)$ denotes the activation output at layer $k$, and $\epsilon$ is a small positive constant to ensure numerical stability (e.g., $\epsilon=10^{-8}$). Minimising this loss preserves compositional integrity across the specified layers by encouraging smooth and consistent transitions between them, thereby enabling more stable information flow and aggregation within this range.
@@ 4.1 Downstream Tasks & datasets
CARMA is evaluated across two tasks that assess different aspects of compositional generalisation: Inverse Dictionary Modelling for word-level composition and Sentiment Classification for phrase-level structure. These tasks measure systematicity, substitutivity, over-generalisation, and robustness to perturbations.
Inverse Dictionary Modelling (IDM) evaluates a model’s ability to generate terms from definitions, focusing on substitutivity in semantic composition. WordNet Miller (1994) is used as the training dataset, with an 80-10-10 train-validation-test split. Models are prompted with a definition and tasked with generating the corresponding term (e.g., “The star around which the Earth orbits is called” $\rightarrow$ “Sun”). Performance is assessed using Exact Match Accuracy, which measures whether the generated term precisely matches the expected output. By mapping definitions to terms, this task provides a robust assessment of a model’s ability to perform compositional substitution.
Sentiment Classification (SC) assesses the model’s ability to infer sentiment from phrases and sentences, particularly focusing on sentiment shifts and over-generalisation. The Stanford Sentiment Treebank (SST) Socher et al. (2013) is used with its original dataset splits. Models predict sentiment labels given textual inputs (e.g., “A brilliant performance sentiment is” $\rightarrow$ “positive”). Performance is evaluated using Exact Match Accuracy. This task examines how sentiment composition is preserved across different levels of linguistic structure. Task formalisation, dataset details, and task selection rationale are in Appendices A, B.1, and B.2, respectively.
4.2 Model Configurations and Baselines
Experiments are conducted across three model configurations: baseline models, models with task-specific fine-tuning, and models with fine-tuning plus CARMA regularisation. Models use 500 warm-up steps and a 0.006 learning rate. We test GPT-2 (S/L) Radford et al. (2019), Gemma-2B Team et al. (2024), Llama (1B/3B) Dubey et al. (2024), and Qwen (0.5B/3B) Yang et al. (2024a), representing diverse architectures and capacities. CARMA regularisation is generally applied at approximately one-third of the model’s depth, though specific layer positions vary. Details on fine-tuning methodologies, model specifications, and CARMA hyperparameter selection are provided in Appendix B.3.
@@ 4.3 Interventions for Compositional Robustness and Performance Stability
Two interventions are used to evaluate the robustness of compositional structures and the stability of learned representations: Constituent-aware pooling and synonym replacement. These interventions assess hierarchical dependencies and semantic consistency under controlled perturbations.
Constituent-Aware Pooling (CAP) Aljaafari et al. (2024) groups token-level representations into higher-level semantic units (e.g., words, syntactic constituents) to assess hierarchical dependencies and how compositional structures are maintained across layers. In this paper, the token-to-word CAP is utilised. Model robustness is measured by monitoring performance metrics before and after applying CAP. Full methodology and formalisation are provided in Appendix C.1.
@@
4.4 Experimental setup
Experiments were conducted using NVIDIA RTX A6000 and A100 GPUs. The method was developed in Python (v3.10.15) with Transformers (v4.44.2) Wolf et al. (2020), PyTorch (v2.4.1) Paszke et al. (2019), and Transformer-lens (v2.8.1) Nanda and Bloom (2022). Preprocessing tasks, including tokenisation and tagging, used NLTK (v3.9.1) Bird et al. (2009), spaCy (v3.7.2) Honnibal et al. (2020), and TextBlob (v0.18.0) Loria et al. , with Scikit-learn (v1.5.1) Pedregosa et al. (2011) for evaluation.
@@ 5 Results and discussion
The method is evaluated across three aspects: (1) its impact on model robustness against compositional-based perturbations, (2) its impact on model performance stability, and (3) its impact on model overall performance. See Appendix B.4 for a detailed breakdown of the evaluation metrics used for each aspect.
@@ 5.1 Constituent-Aware Pooling (CAP) Intervention
Figures LABEL:fig:carma_results_idm and LABEL:fig:carma_results_sst show the impact of CAP on IDM and SC tasks, comparing original, fine-tuned (FT) and CARMA models. [^2] Throughout this paper, models incorporating CARMA with FT are referred to as CARMA models. Model performance is averaged across three CAP protocols (Mean, Max, and Sum), with per-protocol results provided in Appendix E. The analysis examines how well models preserve compositionality under hierarchical pooling.
@@ 5.2 Synonyms Replacement Intervention
Synonym Replacement evaluates semantic consistency and robustness under lexical variations across multiple runs ($N\geq 5$) with different seeds. ConsistSyn measures output preservation after substitution, while the coefficient of variation (CV) quantifies performance stability, with lower values indicating higher stability. Performance is assessed at 25% and 40% word replacement rates to measure sensitivity to increasing perturbations. A sample of results is presented in Table 1, with full details in Appendix E.
@@
Figure 3: Task performance in IDM across GPT-2 (S, L), Gemma-2B, Llama (1B, 3B), and Qwen (0.5B, 3B).
Figure 4: Task performance in SC across GPT-2 (S, L), Gemma-2B, Llama (1B, 3B) and Qwen (0.5B, 3B).
@@ 5.3 Impact of CARMA on Performance
Figures 3 and 4 show the performance of original, FT, and CARMA accuracies across tasks. CARMA demonstrates significant improvements over original models across tasks. For example, in IDM, GPT2-L achieves 150% improvement, and Llama-3B shows an 89.6% increase, while in SC, Gemma-2B demonstrates 122.5% improvement over Original baselines.
Task-specific patterns emerge when comparing models. For instance, in IDM, CARMA outperforms FT, with Llama-3B showing a +5% gain and GPT2-L improving by 1.7%. In SC, CARMA maintains comparable performance to FT while enhancing robustness, suggesting it preserves learned features while strengthening compositional consistency.
@@ 7 Conclusion
This paper presents CARMA, a method for enhancing compositional generalisation in LLMs through mutual information regularisation and stability constraints. By addressing information fragmentation and layer-wise instability, CARMA improves performance stability and robustness under interventions, as demonstrated through IDM and SC tasks. The method offers a cost-effective solution applicable across model architectures with minimal modifications. Future work should explore extending CARMA to additional tasks that rely more on nuanced semantic features and multilingual settings to further evaluate its scalability and adaptability. Integrating CARMA into improved, targeted transformer architectures for CG could unlock new possibilities for enhancing compositionality.
@@
The limitations of this paper can be summed up as follows: First, our results are primarily reported for the English language. Further analysis across languages with diverse linguistic structures is left as a confirmatory future work. Second, the datasets (WordNet and SST) lack a more comprehensive representativeness of broader linguistic phenomena. Third, our focus is predominantly on decoder-based Transformers. Finally, the employed Transformer models may inherit potential biases ingrained from their pre-training data.
@@ B.4 Evaluation Metrics
Synonym Consistency (ConsistSyn) (ConsistSyn)
@@
The coefficient of variation (CV) measures the stability of model performance across multiple runs, with lower values indicating greater consistency. It is defined as:
@@
Normalised Improvement (NI) evaluates the relative gain in consistency introduced by a model over a baseline model. It is calculated as:
@@ C.2 Synonym Replacement
Appendix D InfoNCE for Mutual Information Estimation
Appendix E Extended results
Figures LABEL:fig:idm_models and LABEL:fig:sst_models, and Tables 4 and 5 provide additional results for models’ performance comparison under CAP and synonym interventions. CARMA models show a clear advantage over all models and tasks. However, the gain is clearer in the IDM case, where more intricate features and compositionality generalisation are required. It is also observed that the performance of the FT and CARMA models demonstrates similar curves or trends. Given this observation, we argue that CARMA’s improvements stem from its learning objectives, which align closely with cross-entropy loss while explicitly addressing intermediate representation stability. The observed improvements are moderate in some cases, particularly for SC tasks. This behaviour is expected due to the limited size of the fine-tuning datasets compared to the original pretraining data used for these models. Nevertheless, larger models, such as Llama-3B and Gemma-2B, exhibit more substantial improvements with CARMA, demonstrating its scalability with model capacity.

### v2

@@ Abstract
Nura Aljaafari1,4†,  Danilo S. Carvalho1,3,  André Freitas1,2,3 1 Department of Computer Science, University of Manchester, United Kingdom 2 Idiap Research Institute, Switzerland 3 National Biomarker Centre, CRUK-MI, Univ. of Manchester, United Kingdom {firstname.lastname}@[postgrad.]manchester.ac.uk
@@ 1 Introduction
Despite their strong performance across various NLP tasks, large language models (LLMs) exhibit persistent weaknesses in compositional generalisation Hupkes et al. (2020); Kim and Linzen (2020a); Aljaafari et al. (2024). These limitations stem from multiple factors, including training objectives and model architectures. Standard autoregressive training methods, such as next-token prediction, prioritise statistical correlations in token sequences over structured semantic understanding Yin et al. (2023a); Dziri et al. (2024). As a result, token representations often lack structured compositionality, leading to fragmented information processing within layers (horizontal misalignment) and across layers (vertical inconsistency).
@@ 2 Compositionality in LLMs
Compositional generalisation (CG) in linguistics encompasses five key principles: systematicity, productivity, substitutivity, localism, and over-generalisation Dankers et al. (2022a). These principles have been explored in LLMs across compositional instruction Yang et al. (2024b), semantic parsing Li et al. (2023), translation Li et al. (2021), and multi-step inference Zhang et al. (2024). Studies show standard Transformer-based LLMs exhibit limited CG, struggling with basic compositional tasks such as assembling tokens into words or constructing morphemes Aljaafari et al. (2024); Ismayilzada et al. (2024). These limitations are linked to architectural constraints, training objectives, and tokenisation practices that fragment information and increase sensitivity to input order and contextual noise Murty et al. (2023).
@@
Architectural Mechanisms and Compositional Consistency. Beyond training objectives, architectural mechanisms such as dropout and self-attention disperse information across the model, increasing sensitivity to input order and context. This undermines compositional consistency Sajjadi et al. (2016); Cai et al. (2021), the ability to maintain consistent outputs when processing semantically equivalent inputs through transformations like word substitution or paraphrasing. These challenges impact both high-complexity reasoning tasks and simpler operations that demand consistent morphological and syntactic processing Ismayilzada et al. (2024).
Existing Approaches to Enhance CG in LLMs. Research has explored architectural adjustments, regularisation techniques, and task-specific strategies to address CG limitations. Ontanon et al. (2022) demonstrated that combining relative positional encoding with embeddings enhances CG in algorithmic tasks, while weight sharing and copy decoders help retain input structures. Architectural modifications like Pushdown Layers Murty et al. (2023) and GroCoT Sikarwar et al. (2022) incorporate mechanisms for tracking syntactic depth and spatial relations, enabling recursive processing of compositional structures. RegularGPT Chi et al. (2023) introduces adaptive depth and memory mechanisms to facilitate CG. Studies by Csordás et al. (2021) and Petty et al. (2024) reveal that architectural choices and training setups significantly impact CG enhancement. In neural machine translation, Dankers et al. (2022b) found a positive correlation between data size and compositional performance.
Frameworks like CompMCTG and Meta-MCTG Zhong et al. (2024) suggest joint training and meta-learning approaches improve fluency, though performance drops persist in out-of-distribution tasks. Synthetic tasks show recursive, step-by-step prompt formats support combinatorial generalisation, despite training biases and sequence order constraints Ramesh et al. (2024).
@@ 3 Enhanced Compositionality via Advanced Regularisation and Mutual Information Alignment (CARMA)
This section formalises compositionality, introduces the core principles of CARMA, and details its components. Figure 1 illustrates the method, highlighting its process and key components.
@@ 3.2 CARMA Formalisation
where $f^{(k)}(X)$ denotes the output of layer $k$ (i.e., after the attention and MLP submodules), and $\epsilon$ is a small positive constant to ensure numerical stability (e.g., $\epsilon=10^{-8}$). Minimising this loss preserves compositional integrity across the specified layers by encouraging smooth and consistent transitions between them, thereby enabling more stable information flow and aggregation within this range.
@@
Layer Selection for Regularisation.
We apply CARMA to layers around one-third of the model depth, based on evidence that early-to-mid layers better capture compositional and syntactic structure, while deeper layers tend to specialise in task-specific representations (He et al., 2024; Langedijk et al., 2024). In our preliminary experiments, we observed that performance gains diminish when regularisation is applied to deeper layers. As a default, we recommend layers 3–4 in 12-layer models and 6–10 in 24-layer models.
Evaluation Metrics.
We use exact match accuracy as the primary metric for both IDM and SC. This choice is motivated by the fact that both tasks have closed and categorical output spaces: IDM outputs are limited to a predefined set of single-word lexical entries from WordNet, while SC uses a fixed sentiment label set. In such settings, exact match is a standard and appropriate evaluation criterion.
@@ 4.1 Downstream Tasks & datasets
Two tasks that assess different aspects of compositional generalisation are used in the paper: Inverse Dictionary Modelling (IDM) for word-level composition and Sentiment Classification (SC) for phrase-level structure. These tasks measure systematicity, substitutivity, over-generalisation, and robustness to perturbations.
IDM evaluates a model’s ability to generate terms from definitions, focusing on substitutivity in semantic composition. Using WordNet Miller (1994) with an 80-10-10 train-validation-test split, models are prompted with a definition to generate the corresponding term (e.g., The shore of a sea is called” $\rightarrow$ coast”). By mapping definitions to terms, this task provides a robust assessment of a model’s ability to perform compositional substitution.
SC assesses the model’s ability to infer sentiment from phrases and sentences, particularly focusing on sentiment shifts and over-generalisation. Using the Stanford Sentiment Treebank (SST) Socher et al. (2013) with its original splits, models predict sentiment labels from textual inputs (e.g., A brilliant performance sentiment is” $\rightarrow$ positive”). This task examines how sentiment composition is preserved across different levels of linguistic structure. For both tasks, performance is assessed using Exact Match Accuracy, providing a robust assessment of compositional substitution ability. Task formalisation, dataset details, and task selection rationale are in Appendices A, B.1, and B.2, respectively.
4.2 Models and Experimental setup
We evaluate three setups: original models, task-specific fine-tuning, and fine-tuning with CARMA regularisation. We test GPT-2 (S/L) Radford et al. (2019), Gemma1–2B Team et al. (2024), LLaMA3.2 (1B/3B) Dubey et al. (2024), and Qwen2.5 (0.5B/3B) Yang et al. (2024a). We focus on decoder-only architectures, as they represent the dominant paradigm in many open-weight and production-ready LLMs. CARMA is generally applied at approximately one-third of the model’s depth, though specific layer positions vary. Details on software, FT methodologies, model specifications, and CARMA hyperparameter selection are provided in Appendices B.3 and B.5.
@@ 4.3 Interventions for Compositional Robustness and Performance Stability
Two interventions are used to evaluate the robustness of compositional structures and the stability of learned representations: Constituent-aware pooling (CAP) and synonym replacement. These interventions assess hierarchical dependencies and semantic consistency under controlled perturbations.
CAP Aljaafari et al. (2024) groups token-level representations into higher-level semantic units (e.g., words, constituents) to assess hierarchical dependencies and how compositional structures are maintained across layers. In this paper, the token-to-word CAP is utilised. Model robustness is measured by monitoring performance metrics before and after applying CAP. Full methodology and formalisation are provided in Appendix C.1.
@@ 5 Results and discussion
The method is evaluated across three aspects its impact on: (1) model robustness against compositional-based perturbations, (2) model performance stability, and (3) model overall performance. See Appendix B.4 for a detailed breakdown of the evaluation metrics used for each aspect.
@@ 5.1 Constituent-Aware Pooling (CAP) Intervention
Fig. LABEL:fig:carma_results_idm and LABEL:fig:carma_results_sst show the impact of CAP on both tasks, comparing original, fine-tuned (FT), and CARMA models. [^2] Throughout this paper, models incorporating CARMA with FT are referred to as CARMA models. Each plot shows performance across normalised layer positions, where Accuracy is averaged over three CAP protocols (Mean, Max, Sum); protocol-specific results and extended comparisons are included in Appendix E. The analysis examines how well models preserve compositionality under hierarchical pooling.
@@ 5.2 Synonyms Replacement Intervention
Synonym Replacement evaluates semantic consistency and robustness under lexical variations across multiple runs ($N\geq 5$) with different seeds. ConsistSyn measures output preservation after substitution, while the coefficient of variation (CV) quantifies performance stability, with lower values indicating higher stability. Performance is assessed at 25% and 40% word replacement rates to measure sensitivity to perturbations. Sample results are in Table 1; full details appear in Appendix E.
@@
While the absolute differences in ConsistSyn between CARMA and FT are sometimes modest, particularly at lower replacement rates (e.g., 25%), the stability benefits of CARMA become more evident under increased perturbation (e.g., 40%), where FT models often show degraded consistency. In these higher-variance regimes, CARMA consistently reduces output variability across model families, reinforcing its utility as a robustness-oriented intervention, even when raw accuracy remains comparable.
@@
Figure 3: Task performance in IDM across GPT2 (S, L), Gemma-2B, Llama (1B, 3B), and Qwen (0.5B, 3B).
Figure 4: Task performance in SC across GPT2 (S, L), Gemma-2B, Llama (1B, 3B) and Qwen (0.5B, 3B).
@@ 5.3 Impact of CARMA on Performance
Fig. 3 and 4 show the performance of original, FT, and CARMA accuracies across tasks. CARMA demonstrates significant improvements over original models across tasks. For example, in IDM, GPT2-L achieves 150% improvement, and Llama-3B shows an 89.6% increase, while in SC, Gemma-2B demonstrates 122.5% improvement over Original baselines.
Task-specific patterns emerge when comparing models. For example, in IDM, CARMA outperforms FT, with Llama-3B showing a +5% gain and GPT2-L improving by 1.7%. In SC, CARMA maintains comparable performance to FT while enhancing robustness, suggesting it preserves learned features while strengthening compositional consistency.
@@ 7 Conclusion
This paper presents CARMA, a method for enhancing compositional generalisation in LLMs through mutual information regularisation and layer-wise stability constraints. By addressing information fragmentation and instability across layers, CARMA improves performance robustness and stability under intervention. The method requires no architectural changes and integrates cleanly into standard fine-tuning pipelines. Future work should explore extending CARMA to tasks requiring more nuanced semantic reasoning and to multilingual contexts. Another important direction is combining CARMA with techniques that explicitly challenge generalisation, such as adversarial perturbations or structured distribution shifts, to promote the acquisition of novel compositional behaviours. Incorporating CARMA into improved, task-targeted architectures may further enhance its effectiveness. Additionally, controlled training-from-scratch studies could isolate CARMA’s impact more precisely and reveal deeper insights into how it shapes compositional representations across training regimes.
@@
The limitations of this paper can be summed up as follows: First, our results are primarily reported for the English language. Further analysis across languages with diverse linguistic structures is left as a confirmatory future work. Second, the datasets (WordNet and SST) lack a more comprehensive representativeness of broader linguistic phenomena. Third, our focus is predominantly on decoder-based Transformers, and the employed Transformer models may inherit potential biases ingrained from their pre-training data. Finally, while CARMA maintains inference efficiency, it introduces training-time overhead due to auxiliary losses, which should be considered when integrating the method into resource-constrained environments.
@@ B.4 Evaluation Metrics
Synonym Consistency (ConsistSyn)
@@
measures the stability of model performance across multiple runs, with lower values indicating greater consistency. It is defined as:
@@
evaluates the relative gain in consistency introduced by a model over a baseline model. It is calculated as:
@@
B.5 Experimental setup
Experiments were conducted using NVIDIA RTX A6000 and A100 GPUs. The method was developed in Python (v3.10.15) with Transformers (v4.44.2) Wolf et al. (2020), PyTorch (v2.4.1) Paszke et al. (2019), and Transformer-lens (v2.8.1) Nanda and Bloom (2022). Preprocessing tasks, including tokenisation and tagging, used NLTK (v3.9.1) Bird et al. (2009), spaCy (v3.7.2) Honnibal et al. (2020), and TextBlob (v0.18.0) Loria et al. , with Scikit-learn (v1.5.1) Pedregosa et al. (2011) for evaluation. Models use 500 warm-up steps and a 0.006 learning rate.
@@ C.2 Synonym Replacement
Appendix D InfoNCE for Mutual Information Estimation
Appendix E Extended results
Figures LABEL:fig:idm_models, LABEL:fig:sst_models and Tables 4 and 5 provide additional results for models’ performance comparison under CAP and synonym interventions. Figure LABEL:fig:carma_results_combined_full shows an overall models performance under cap for all models. CARMA models show a clear advantage over all models and tasks. However, the gain is clearer in the IDM case, where more intricate features and compositionality generalisation are required. It is also observed that the performance of the FT and CARMA models demonstrates similar curves or trends. Given this observation, we argue that CARMA’s improvements stem from its learning objectives, which align closely with cross-entropy loss while explicitly addressing intermediate representation stability. The observed improvements are moderate in some cases, particularly for SC tasks. This behaviour is expected due to the limited size of the fine-tuning datasets compared to the original pretraining data used for these models. Nevertheless, larger models, such as Llama-3B and Gemma-2B, exhibit more substantial improvements with CARMA, demonstrating its scalability with model capacity.
E.1 Training Runtime and Overhead
We report wall-clock training times (in minutes) for each model under standard fine-tuning (FT) and with CARMA. Overhead is computed as the relative increase in runtime caused by the additional mutual information and stability losses. All experiments for models were conducted on a single GPU under identical batch size, optimiser settings, and hardware configuration.
Model FT (min) CARMA (min) Overhead ratio
GPT2 Small 1.9 3.48 ×$\times 1.8$
GPT2 Large 6.7 8.9 ×$\times 1.3$
Llama 3.2–1B 2.78 6.96 ×$\times 2.5$
Llama 3.2–3B 7.55 20.2 ×$\times 2.2$
Qwen 2.5–0.5B 2.10 4.75 ×$\times 2.3$
Qwen 2.5–3B 2.98 5.01 ×$\times 1.7$
Table 6: Wall-clock IDM training time and overhead introduced by CARMA. All runs use a single GPU under identical batch, optimiser, and hardware settings.
Model FT (min) CARMA (min) Overhead ratio
GPT-2 Small 1.12 8.51 ×$\times 7.6$
GPT-2 Large 6.02 13.18 ×$\times 2.19$
LLaMA 3.2–1B 2.50 9.98 ×$\times 4.9$
LLaMA 3.2–3B 6.58 14.06 ×$\times 2.13$
Qwen 2.5–3B 6.01 16.86 ×$\times 2.8$
Table 7: Wall-clock SC training time and overhead introduced by CARMA. All runs use a single GPU under identical batch, optimiser, and hardware settings.
CARMA introduces non-trivial training-time overhead due to auxiliary objectives, particularly for smaller models or longer sequences. However, inference costs remain unchanged, and no architectural modifications are required. We observe moderate to high training slowdowns (e.g., $\times$2.2–$\times$2.9 for LLaMA-3B), with variance across models due to token length and loss computation. Optimising runtime for mutual information and stability estimation is an important direction for future efficiency improvements.